OpenStudy (anonymous):

can someone graph this for me or tell me the points for: 2x+3y=8 -x+y=-4

OpenStudy (anonymous):

(4,0)

OpenStudy (anonymous):

Set x= 0 to get y and vice versa for each equation. Join the points up.

OpenStudy (anonymous):

2x+3y=8 -x+ y=-4 Multiply by two 2x+3y=8 -2x+2y=-8 Cancel the x's (2x-2x), add the y's and the integers 5y = 0 therefore y = 0 if y = 0 then: 2x+3(0) = 8 2x = 8 x = 4, y = 0 (4,0)

OpenStudy (anonymous):

U want to graph it, right?

OpenStudy (anonymous):

thansk what about this one x-2y=4 4x+2y=6

OpenStudy (anonymous):

Try the same thing as I did before, just with the new equation. Post the answer you get.